OpenStudy (anonymous):

If the ratio of three sides of a triangle is 1/4:1/8:1/6 and the perimeter is 4.75ft, what is the length of the longest side?

OpenStudy (anonymous):

0.25+1/8+1/6 = 13/24 4.75/(13/24)=114/13 multiply it with 1/4 to find longest side: (114/13)*(1/4)=57/26 i really hope this helps

OpenStudy (anonymous):

problem here are the numbers, especially 1/6, which is making the problem here, this would've been lot easier if it was some different number

OpenStudy (anonymous):

Yeah, this helped a ton! Thanks so much! If it weren't for that 1/6 I probably would have gotten it, but I also missed the lesson because I was sick, just found the homework on the website and I have a copy of the textbook at home so might as well.